# Design of a Quality Management System based on the EU Artificial Intelligence Act

### v1

@@ 5 Conclusion
The presented quality management system (QMS) is a tool for ensuring compliance with the EU AIA regulations for high-risk AI and GPAI systems, guided by legal and system design functional (FR) and non-functional requirements (NFR). The QMS is based on a microservice architecture, directly connecting AI systems and containing several sub-services, each with a different purpose based on the regulations for high-risk AI or GPAI systems of the EU AIA. This SaaS web application aims to map the compliance management processes for AI systems (especially high-risk AI and GPAI systems) into one tool and to carry them out efficiently. Quantitative tests can be directly applied to the adopted AI system within the QMS. Although the presented first version prototype QMS is optimized to check and document LLMs, its concept, design, and architecture can be applied to various types of AI systems. Currently, the QMS integrates a risk management system (RMS) (cf. Article 9 EU AIA) and a data management and governance system (DMDGS) (cf. Article 10 EU AIA) as sub-services. Further sub-services, such as logging the use of the AI system (Article 12 EU AIA), transparency approaches on how to create instructions for use (cf. Article 13 EU AIA), and a human-machine interface for users of the AI system (cf. Article 14 EU AIA), will be added in future versions of the prototype QMS. Additionally, each sub-service sub-component can be improved in the future, particularly the identification and mitigation components requiring stakeholder and AI user involvement. Currently, risk identification is not linked with risk analysis, such as by suggesting technical evaluation metrics or tools for addressing potential risks manually or automatically. Future work will focus on better integrating risk identification results with the risk analysis and assessment components. The idea of this paper is related to previous compliance management work (cf. [22, 23]). The goal for future versions of the prototype QMS is to integrate findings and technical features from earlier automated compliance verification research and insights from business process management research, such as modeling larger processes by combining regulations for AI development and post-development when the system is in use.

### v2

@@ 5 Conclusion
The presented QMS is a tool for ensuring compliance with the AIA regulations for high-risk AI and GPAI systems, guided by legal and system design requirements. The QMS is based on a microservice architecture, directly connecting AI systems and containing several sub-services, each with a different purpose based on the regulations of the AIA. This SaaS web application aims to map the compliance management processes for AI systems into one tool and to carry them out efficiently. Assessments can be directly applied to the adopted AI system within the QMS. Although the first version of the presented prototype QMS is optimized to check and document LLMs, its concept, design, and architecture can be applied to various AI systems. Currently, the QMS integrates a risk management system and a data management and governance system as sub-services. Further sub-services, such as logging the use of the AI system (Art. 12 AIA), transparency approaches on how to create instructions for use (cf. Art. 13 AIA), and a human-machine interface for users of the AI system (cf. Art. 14 AIA), will be added in future versions of the prototype QMS. Future versions of the prototype QMS aim to integrate findings and technical features from earlier automated compliance verification research and insights from business process management research, such as modeling more extensive processes by combining regulations for AI development and post-development when the system is in use. Additionally, future work will further develop the various sub-services and conduct a broader evaluation to make the QMS even more comprehensive.